A small piece of dust has a net negative charge of -2.1 micro coulombs. How many extra electrons does it have?
Answer #1

charge on 1 electron = -1.6*10^-19 C

charge on the piece of dust = -2.1*10^-6

number of excess electrons = (-2.1*10^-6 / -1.6*10^-19)

= 1.3125*10^13